The sum of two numbers is 7. One of the numbers is five less than the other. Find the numbers.
Enter your answer as a list of numbers separated by a comma: a,b

Found 2 solutions by mananth, josgarithmetic:
Answer by mananth(16946)

The sum of two numbers is 7. One of the numbers is five less than the other. Find the numbers.
Enter your answer as a list of numbers separated by a comma: a,b
Let one number be a
other be b
a+b=7.................................................1
assume a is bigger number
a-b=5.............................................2
add 1 & 2
we get 2a=12
a=6
therefore b is 1
write the answer in the form asked for

Answer by josgarithmetic(39617)

n and n-5
sum is 7

..
2n-5=7
n=6
Second time THIS example was answered!